ABC, Inc. has a debt-equity ratio of 80%. The firm is analyzing a new project which requires an initial cash outlay of $300,000 for new equipment. The flotation cost for new equity is 9 % and for debt 4.5 %. (a) What is the weighted average flotation cost? (b) What is the initial cost of the project including the flotation costs?
Answer #1

a. Assume equity is 1

Debt equity ratio = Debt / Equity

0.8 = Debt / 1

Debt = 0.8

Equity = 1

Debt + Equity = 1.8

Weighted average floatation cost = (9% *1/ 1.8) +(4.5% * 0.8/ 1.8) = 7%

b. Initial cost = 300,000/ (1-7%) = 322,581 (rounded off)